# Trust, but verify

### v1

@@ 1 Introduction
Two years after OpenAI released ChatGPT, there is an explosion of open-source LLMs both in quality and quantity [Zhao et al. 2025]. The best open-source LLMs are now matching the best close-source LLMs in performance [Fourrier, Habib, Lozovskaya, Szafer & Wolf 2024], while offering much more variety in terms of model size, speed, and specific "skills", such as math and coding [Jiang, Wang, Shen, Kim & Kim 2024]. The rise of open source enables individual users to run customized LLMs and agent applications on their own computers without centralized gatekeepers. The decentralized LLM inference has many benefits, such as privacy, low cost, speed, and availability. But, more importantly, it fosters an ecosystem of customized LLM services based on models fine-tuned and supplemented by proprietary data and knowledge.
Decentralized AI inference networks, such as Gaia [Gaia Foundation 2024], allow individuals to pool together their computing resources to serve the in-demand customized LLMs to the public and share revenues with the model creators. Each Gaia node runs a set of AI models, including LLMs, and a knowledge base. It advertises and makes available its services to the public. However, as those networks must stay permissionless, to reduce censorship, protect privacy and lower the barrier of participation, it is possible for some Gaia nodes to cheat and run a different LLM or knowledge base than the ones it claims to run. A popular Gaia "domain", which is a sub-net of nodes that are supposed to run the same LLM and knowledge base, could have over 1,000 nodes at any given time. The network must have a mechanism to automatically detect and penalize the nodes that cheat.
@@ 2 Past research
Zero Knowledge Proof (ZKP) allows us to verify the outcome of a computation process without knowing the internals of the computing system [Goldwasser, Micali & Rackoff 1985]. In the context of LLM verification, we can build a ZKP circuit for an LLM, which represents the inference process. The LLM service provider (Gaia node) generates a proof for each incoming request (i.e., the LLM prompt) indicating that it has correctly executed the inference process [Sun, Li & Zhang 2024]. Any external party (validator) can then check the proof and confirm that it is generated by the specific LLM without knowing the exact software setup on the machine.
@@
  • Even the state-of-the-art ZKP algorithms running on expensive Nvidia GPUs require 13 minutes to generate proof for a single inference execution for a small 13B LLM [Sun, Li & Zhang 2024]. That is 2 orders of magnitude slower (and 2 orders of magnitude more resource consuming) than the inference task itself. Since LLM inference is already compute-bound, spending 100×100$\times$100 × on compute just for verification is not feasible.
  • The memory cost of generating ZKP proofs is also staggering. A toy LLM with one million parameters would require over 25GB of RAM even with the most optimized algorithms [Ganescu & Passerat-Palmbach 2024]. Scaling that to LLMs of 10,000×10,000$\times$10 , 000 × of that size would require computing resources far beyond any reasonable computer.
@@
  • The TEE reduces raw hardware (CPU) performance by a factor of up to 2×2$\times$2 × [Dong & Wang 2025], which is very hard to accept since LLM inference is already a compute-bound task.
  • There are very few GPUs or AI accelerators that support TEEs. The high-end H-100 is the first Nvidia GPU with TEE support [Apsey, Rogers, O’Connor & Nertney 2023].
@@
  • Distributing private keys to decentralized and permissionless TEE devices is a major challenge that requires specialized software and operational procedures [Oasis 2025].
@@
EigenLayer’s AVS [EigenLayer 2023] provides a set of smart contracts that allows independent operators and validators to stake crypto assets, and vote on intersubjective results from the sampling. Once a vote reaches consensus, known as an intersubjective consensus [EigenLayer 2024], the smart contracts automatically execute to reward or punish the operators and validators.
@@ 3 Hypotheses
Recall that a Gaia node is a network server that runs a set of AI models and knowledge bases. It promises to run the AI and knowledge base stack it advertises on the Gaia network.
@@ 4 Experiments
Hence we ran two sets of experiments to test the above hypotheses on LLM and knowledge base variations respectively. The source code and data to replicate all the analyses and results in this paper are available as an open source GitHub repo [Gaia Foundation 2025a].
@@ 4.1 Responses from different models
  • Llama 3.1 8b by Meta AI [Grattafiori et al. 2024].
  • Gemma 2 9b by Google [Mesnard et al. 2024].
  • Gemma 2 27b by Google [Mesnard et al. 2024].
@@ 4.2 Responses from different knowledge bases
We set up two Gaia nodes with the same LLM, Gemma-2-9b by Google [Mesnard et al. 2024]. The knowledge bases are vector databases created from text files [Gaia Foundation 2025b]. In this experiment, we adopted knowledge bases from Paris and London Wikipedia pages respectively [Wikipedia 2025a, 2025b]. The knowledge bases are created with the nomic-embed-text-1.5 embedding model [Nussbaum, Morris, Duderstadt, and Mulyar, 2025]. The vector database files are available as follows.
  • Paris [Yuan 2024a].
  • London [Yuan 2024b].
@@ 5 Results
  • Generated an embedding vector using the gte-Qwen2-1.5B-instruct embedding model [Yang et al. 2024]
@@ 5.1 Distinguishing different models
Critically, Table 1 shows that the distances between model pairs are $32-65$ times larger than the highest RMS scatter observed within any model ($0.0062$ for Llama-3.1-8b). This large separation between inter-model distances and intra-model variation indicates that Gaia nodes with different LLMs produce reliably distinguishable outputs, making them identifiable through their response patterns.
@@
Gemma9b - Gemma27b $0.1558$ 32.5×32.5$\times$32.5 ×
Llama8b - Gemma9b $0.3129$ 65.2×65.2$\times$65.2 ×
Llama8b - Gemma27b $0.3141$ 65.4×65.4$\times$65.4 ×
@@ 5.2 Distinguishing different knowledge bases
Critically, the distances between knowledge base pairs (average $0.0862$) are about $5-26$ times larger than the RMS scatter observed within a knowledge base ($0.0072$). This large separation between inter-knowledge base distances and intra-knowledge base variation indicates that Gaia nodes with different knowledge bases produce reliably distinguishable outputs, making them identifiable through their response patterns.
@@ 6 Discussions
The two Gemma models are more similar to each other (distance of $0.1558$) than either is to Llama (distances around $0.31$). While still 32×32$\times$32 × greater than the observed internal variation, this suggests that distinguishing between models within the same family might present different challenges than distinguishing between models from different families.
@@
Finally, different questions demonstrated varying levels of differentiation between models or knowledge bases. For example, Q13 ("Who wrote ’Romeo and Juliet’?") showed the greatest distance between models ($0.5291$ between Llama 8b and Gemma 27b). Q18 ("How many bridges did Philip Augustus build in Paris in the late 12th century?") showed the greatest distance between knowledge bases ($0.188$ between London and Paris). Those variations in the effectiveness of test questions are important in practical applications.
@@ 7 System design and implementation
In this paper, we propose an AVS design that is tailored to the Gaia network architecture. The AVS is organized into multiple "operator sets".
@@
  • Poll each Gaia node in a domain using a random question from a question set that is specially designed for this domain. The question set will be provided by the domain owner.
@@
  • Post the results on EigenDA [EigenDA 2025].
At the end of each Epoch, a randomly chosen validator will act as the aggregator and perform a map-reduce function on all messages posted by all validators in that Epoch. The result is a set of flags for each Gaia node, such as

### v2

@@ 1 Introduction
Two years after OpenAI released ChatGPT, there is an explosion of open-source LLMs both in quality and quantity [1]. The best open-source LLMs are now matching the best closed-source LLMs in performance [2], while offering much more variety in terms of model size, speed, and specific "skills", such as math and coding [3]. The rise of open source enables individual users to run customized LLMs and agent applications on their own computers without centralized gatekeepers. The decentralized LLM inference has many benefits, such as privacy, low cost, speed, and availability. But, more importantly, it fosters an ecosystem of customized LLM services based on models fine-tuned and supplemented by proprietary data and knowledge.
Decentralized AI inference networks, such as Gaia [4], allow individuals (and businesses) to run AI models on their own computers. Each "Gaia node" runs a set of AI models, including LLMs, and a knowledge base. It advertises and makes available its services to the public.
Gaia nodes can pool together their computing resources into a "Gaia domain", which is a set of Gaia nodes that are supposed to run the same LLM and knowledge base. A Gaia domain load-balances across its Gaia nodes to serve the in-demand customized LLMs to the public and share revenues with the model creators. A popular Gaia domain could have over 1,000 nodes at any given time. However, as those networks must stay permissionless, to reduce censorship, protect privacy and lower the barrier of participation, it is possible for some Gaia nodes to cheat and run a different LLM or knowledge base than the ones it’s domain requires. The network must have a mechanism to automatically detect and penalize the nodes that cheat.
@@ 2 Past research
Zero Knowledge Proof (ZKP) allows us to verify the outcome of a computation process without knowing the internals of the computing system [5]. In the context of LLM verification, we can build a ZKP circuit for an LLM, which represents the inference process. The LLM service provider (Gaia node) generates a proof for each incoming request (i.e., the LLM prompt) indicating that it has correctly executed the inference process [6]. Any external party (validator) can then check the proof and confirm that it is generated by the specific LLM without knowing the exact software setup on the machine.
@@
  • Even the state-of-the-art ZKP algorithms running on expensive Nvidia GPUs require 13 minutes to generate proof for a single inference execution for a small 13B LLM [6]. That is 2 orders of magnitude slower (and 2 orders of magnitude more resource consuming) than the inference task itself. Since LLM inference is already compute-bound, spending $100\times$ on compute just for verification is not feasible.
  • The memory cost of generating ZKP proofs is also staggering. A toy LLM with one million parameters would require over 25GB of RAM even with the most optimized algorithms [7]. Scaling that to LLMs of $10,000\times$ of that size would require computing resources far beyond any reasonable computer.
@@
  • The TEE reduces raw hardware (CPU) performance by a factor of up to $2\times$ [8], which is very hard to accept since LLM inference is already a compute-bound task.
  • There are very few GPUs or AI accelerators that support TEEs. The high-end H-100 is the first Nvidia GPU with TEE support [9].
@@
  • Distributing private keys to decentralized and permissionless TEE devices is a major challenge that requires specialized software and operational procedures [10].
@@
The validators could compare LLM outputs in the embedding space [11]. Each word token is mapped to a vector of real numbers in a high-dimensional space (e.g., 768 ot 1536 dimensions). The vector, known as an embedding, represents the initial (or most common) meaning of the word token. The embeddings in a sentence or paragraph are then sequentially transformed by a specially trained LLM to adjust for its context (e.g., the tokens before it) in the sentence. The last token has supposedly "seen" every word in the sentence, and hence its context-adjusted embedding is taken as the overall embedding of the sentence. The result is that sentences that have similar meanings will have embedding vectors that are close together. That allows the validators to compare the similarity of outputs from Gaia nodes.
EigenLayer’s AVS [12] (Actively Validated Service or Autonomous Verifiable Service) provides a set of smart contracts that allows independent operators and validators to stake crypto assets, and vote on intersubjective results from the sampling. Once a vote reaches consensus, known as an intersubjective consensus [13], the smart contracts automatically execute to reward or punish the operators and validators.
@@ 3 Hypotheses
Recall that a Gaia node is a network server that runs a set of AI models and knowledge bases. It promises to run the AI model and the knowledge base it advertises on the Gaia network.
@@ 4 Experiments
Hence we ran two sets of experiments to test the above hypotheses on LLM and knowledge base variations respectively. The source code and data to replicate all the analyses and results in this paper are available as an open source GitHub repo [14].
@@ 4.1 Responses from different models
  • Llama 3.1 8b by Meta AI [15].
  • Gemma 2 9b by Google [16].
  • Gemma 2 27b by Google [16].
@@ 4.2 Responses from different knowledge bases
We set up two Gaia nodes with the same LLM, Gemma-2-9b by Google [16]. The knowledge bases are vector databases created from text files [17]. In this experiment, we adopted knowledge bases from Paris and London Wikipedia pages respectively [18, 19]. The knowledge bases are created with the nomic-embed-text-1.5 embedding model [20]. The vector database files are available as follows.
  • Paris [21].
  • London [22].
@@ 5 Results
  • Generated an embedding vector using the gte-Qwen2-1.5B-instruct embedding model [23]
@@ 5.1 Distinguishing different models
Critically, Table 1 shows that the distances between model pairs are 326532-65 times larger than the highest RMS scatter observed within any model ($0.0062$ for Llama-3.1-8b). This large separation between inter-model distances and intra-model variation indicates that Gaia nodes with different LLMs produce reliably distinguishable outputs, making them identifiable through their response patterns.
@@
Gemma9b - Gemma27b $0.1558$ $32.5\times$
Llama8b - Gemma9b $0.3129$ $65.2\times$
Llama8b - Gemma27b $0.3141$ $65.4\times$
@@ 5.2 Distinguishing different knowledge bases
Critically, the distances between knowledge base pairs (average $0.0862$) are about 5265-26 times larger than the RMS scatter observed within a knowledge base ($0.0072$). This large separation between inter-knowledge base distances and intra-knowledge base variation indicates that Gaia nodes with different knowledge bases produce reliably distinguishable outputs, making them identifiable through their response patterns.
@@ 6 Discussions
The two Gemma models are more similar to each other (distance of $0.1558$) than either is to Llama (distances around $0.31$). While still $32\times$ greater than the observed internal variation, this suggests that distinguishing between models within the same family might present different challenges than distinguishing between models from different families.
@@
Finally, different questions demonstrated varying levels of differentiation between models or knowledge bases. For example, Q13 ("Who wrote ’Romeo and Juliet’?") showed the greatest distance between models ($0.5291$ between Llama 8b and Gemma 27b). Q18 ("How many bridges did Philip Augustus build in Paris in the late 12th century?") showed the greatest distance between knowledge bases ($0.188$ between London and Paris). Those variations in the effectiveness of test questions are important in practical applications. We will design a system that can identify and rank the most effective test questions through real-world use cases.
@@ 7 System design and implementation
In this paper, we propose an EigenLayer AVS (Actively Verified Services) design that is tailored to the Gaia network architecture. The AVS is organized into multiple "operator sets".
@@
  • Poll each Gaia node in a domain using a random question from a question set that is specially designed for this domain.
@@
  • Post the results on EigenDA [24].
The initial question set will be provided by the domain owner. But as the AVS runs over time, the validators will collect and filter real-world user questions that are most effective in generating tightly clustered answers from the majority of operator nodes.
At the end of each Epoch, a randomly chosen validator will act as the aggregator and perform a map-reduce function on all messages posted by all validators in that Epoch. For applications in regulated industries, it is possible to utilize Secure Multi-Party Computation (SMPC) methods so that confidential data is not disclosed to validators and the aggregator. The result is a set of flags for each Gaia node, such as
@@
8 Cryptoeconomic considerations
The AVS validators detect outlier nodes that provide "abnormal" answers to the test questions. The approach is statistical and non-deterministic. There could be errors in both type I (false positive, in which an honest node is detected incorrectly) and type II (false negative, in which an dishonest node is not detected). Hence, the "punishment" of supposedly "dishonest" nodes should be gradual and reversible. For exmaple,
  • We should have strict slashing rules and conditions. First-time outliers should not be slashed. The AVS could suspend them so that they do not receive any revenue from AI applications and do not receive staking awards.
  • Repeated outliers and offenders could get slashed. But depending on the severity of the violation, the validator could slash the partial or the full amount of the stake.
The AVS validators are also staked. If a validator consistently disagrees with the majority of validators, he could also be suspended or even slashed.
It is possible for the cryptoeconomic model to fail if the AVS staked value becomes too low. In that case, the outlier nodes would receive very light "punishment" for their dishonest behavior. On the other hand, requiring a large stake would increase the barrier of entry for node operators, which makes the system less secure and less capable by reducing the diversity across node operators.
It is also possible for Gaia nodes in a domain to collude. They could all run an AI model that is different from the domain’s specification. In that case, the AVS will force all nodes in the domain to run the majority-chosen AI model. The AVS will remain self-consistent but it could mis-match the domain’s public advertisement or declaration. Users will find poor service quality from that Gaia domain, and will switch to alternative Gaia domains. The free market will ensure that the honest and healthy domains (i.e., free of collusion) will capture the most economic value.
Each AVS operator set (i.e., Gaia domain) can only scale to 1000 nodes at the time of this writing. There is no inherent limit on how many operators a Gaia validator can check. The Gaia AVS validator could broadcast the test question to all its operator nodes in one go and receive responses asynchronously. The 1000-node limit is a technical limitation for the validator nodes to form consensus in Eigen’s suite of smart contracts. As Eigen improves the efficiency of its system, we expect that the limit will be drastically increased or even removed.
Finally, other AVS approaches have been suggested to support verifiable agents through on-chain smart contracts. However, the current computational limits for on-chain smart contracts make it impossible to run LLM inference on-chain. Eigen’s Layer 1 Agents proposal [25] focuses on verifiable LLM tools, such as function calls and memory use. It calls for LLM tools to be hosted on their own AVSes so that we can ensure that LLMs are getting verifiable data and performing verifiable actions. Those AVSes for LLM tools are highly complementary to the LLM inference AVSes discussed in this paper.